If $6, 700 is invested at a nominal interest rate of 6% per year, compounded semiannually, find the value of the investment after 15 years.
Answer #1

Information provided:

Present value= $6,700

Time= 15 years*2= 30 semi-annual periods

Interest rate= 6%/2= 3% per semi-annual period

The question is solved by computing the future value.

Enter the below in a financial calculator to compute the future value:

PV= -6,700

N= 30

I/Y= 3

Press the CPT key and FV to compute the future value.

The value obtained is 16,262.6586

Therefore, the value of the investment after 15 years is $16,262.66.

The future value can also be calculated using the below formula:

FV=PV (1+r)^t

where:

FV=Future value

PV=Present value

r=rate of interest

t=time period

FV= 6,700*(1+0.03)^30

     = 6,700*2.4273

     = $16,262.66.

In case of any query, kindly comment on the solution